# Non-Clifford and parallelizable fault-tolerant logical gates on constant and almost-constant rate homological quantum LDPC codes via higher symmetries

### v2

@@ III.3 Logical CZ gates as 1-form symmetries
$U^{\prime}_{i,j}(\alpha_{2})=$ $(-1)^{\int_{\mathcal{L}}(\hat{m}\beta^{1})\cup(\hat{l}\gamma^{1})\cup\alpha^{1}}=\left[(-1)^{\hat{m}\hat{l}}\right]^{\int_{\mathcal{L}}\alpha^{1}\cup\beta^{1}\cup\gamma^{1}}$
$=$ $\overline{\text{CZ}}[(\beta^{1};i),(\gamma^{1};j)]^{\int_{\mathcal{L}}{\alpha^{1}\cup\beta^{1}\cup\gamma^{1}}}$
$=$ $\overline{\text{CZ}}[(\beta_{2};i),(\gamma_{2};j)]^{{|\alpha_{2}\cup\beta_{2}\cup\gamma_{2}|}},$ (46)
@@ IV.1 Review of hyperbolic geometry and codes
For $D$$=$$2$, i.e., 2D arithemetic hyperbolic manifold $\mathcal{M}^{2}_{h}$, we can obtain the lower bound of the 1-systole
@@ IV.2 Betti number and encoding rate scaling in three and higher dimensions
where $\chi(\mathcal{M}^{2m})$ is the Euler characteristics and $Pf(\Omega)$ is the Pfaffian of the curvature form of the Levi-Civita connection. Since the Euler characteristics can be expressed as an alternating sum of the $i^{\text{th}}$ Betti number as $\chi(\mathcal{M}^{2m})=\sum_{i=0}^{2m}(-1)^{i}b_{i}$, it is in general difficult to obtain a simple scaling between the $i^{\text{th}}$ Betti number $b_{i}$ and the hyperbolic volume in dimension higher than two from this formula. The only exception is in 4D, where the 2nd Betti number is lower bounded by the volume in Ref.[54], leading to the corresponding $(2,2)$-4D hyperbolic code 888(2,2) means both the logical-X and logical-Z operators are supported on 2-cycles. having constant encoding rate similar to the case of 2D hyperbolic code. Furthermore, in odd dimensions, the Chern-Gauss-Bonnet theorem simply does not apply.
@@
in a quite simple way from an information-theoretic perspective. For a given $D$-manifold, one can define $D-1$ copies of homological quantum codes (with $\lfloor D/2\rfloor$ distinct types due to Poincaré duality), where the physical qubits are assigned to the $i$-cells and the code spaces are associated with $\mathcal{H}_{(i)}=\mathbb{C}^{|H_{i}(\mathcal{M}^{D};\mathbb{Z}_{2})|}$ repsectively ($i=1,2,\cdots,D-1$). Consider the case that $\mathcal{M}^{D}$ only has one connected component, we have $b_{0}(\mathcal{M}^{D})=b_{D}(\mathcal{M}^{D})=1$. Therefore $\sum_{i=0}^{D}b_{i}(\mathcal{M}^{D})-2$ is just the total number of logical qubits in these $D-1$ copies of homological quantum codes. On the other hand, $vol(\mathcal{M}^{D})$ is just proportional to the total number of physical qubits. Therefore, we see that the asymptotic scaling of the Betti number has to be at most linear with the volume as suggested in Eq. (55), simply because one cannot encode more logical qubits than the number of physical qubits. Namely, a Betti number scaling like $O(vol(\mathcal{M}^{D})^{1+\epsilon})$ for any positive $\epsilon$ is impossible.
@@
Figure 8: (a) Illustration of the 3-torus $T^{3}$ as a product of a 2-torus $T^{2}$ and a circle $S^{1}$. (2) Three torus submanifold (2-cycles) $a\times c$, $b\times c$ and $a\times b$ have a triple intersection point at $p$. (c) Illustration of the product manifold $\Sigma_{g}\times S^{1}$. (d) Submanifolds (2-cycles) $\Sigma_{g}$, $a^{(i)}\times c$ and $b^{(i)}\times c$ have a triple-intersection at point $p^{(i)}$. The total triple intersection number is $g$.
@@ V.1.2 Triple intersection and logical gates
There exists a family of quasi-hyperbolic color code with encoding rate $O(1/\log(n))$ and distance $d$$=$$O(\log(n))$, which supports fault-tolerant logical non-Clifford gates and parallelizable logical Clifford gates.
@@ V.2.1 Code construction and code parameters
We first review the construction of such 3-manifolds. We start with the choice of the fibre being a particular family of arithmetic hyperbolic 2-manifolds $\Sigma_{g}$ which are the quotient of the infinite hyperbolic plane $\mathbb{H}^{2}$ by a Fuchian group. We consider the co-compact torsion free Fuchscian group $\Gamma_{(-1,p)}$ derived from the type $\frac{-1,p}{\mathbb{Q}}$ quaternion algebra with the prime number $p=3\ \text{mod}\ 4$. One can explicitly write out the group $\Gamma_{(-1,p)}$ as
@@
    The map $\Sigma_{g}\rightarrow\Sigma_{g}/\langle\tau\rangle$ is a covering projection to the base surface ${}_{g}S=:\Sigma_{g}/\langle\tau\rangle$ (of genus $<g$), which has injectivity radius $R(_{g}S)\geq c_{3}(\log g)^{\frac{1}{2}}$.
where $c_{i}$ ($i$=1,2,3) are positive constants independent of $g$. We now demonstrate the above properties in the following paragraphs.
@@
since both surfaces are obtained from compactifying $\mathbb{H}^{2}$ by quotienting out the corresponding isometry groups. Combined with the Gauss-Bonnet theorem Eq. (48), we know that the genus $g(N)\equiv genus(\mathbb{H}^{2}/\Gamma_{(-1,p)}(N))$ is proportional to the index $[\Gamma_{(-1,p)}:\Gamma_{(-1,p)}(N)]$. Using Eq. (74), we obtain
@@
according to Eq. (78). Choose a base point $q\in\Sigma_{g_{k}}$ for $\alpha$, and then lift $\alpha$ to $\tilde{\alpha}$ in $\Sigma_{g_{k}}$. We now introduce the isometry $\tau_{k}:\Sigma_{g_{k}}\rightarrow\Sigma_{g_{k}}$, satisfying $\tau_{k}\in\Gamma_{(-1,p)}(J_{k})$. The isometry $\tau_{k}$ connects both ends of $\tilde{\alpha}$. The concatenated curve $\tilde{\alpha}(\tau_{k}\tilde{\alpha})(\tau_{k}^{2}\tilde{\alpha})\cdots(\tau_{k}^{n-1}\tilde{\alpha})$ is a closed geodisic loop, which is non-contractible and hence to be longer than the injectivity radius of $\Sigma_{g_{k}}$. This implies that
@@
In the above, we have demonstrated the properties of the hyperbolic surfaces $\Sigma_{g}$. We now construct a fibre bundle with $\Sigma_{g}$$=$$\mathbb{H}^{2}/\Gamma_{(-1,p)}(N)$ as the fibre and a unit-length circle $S^{1}\equiv[0,1]$ as the basis, which is also called a mapping torus construction in this special case (see Sec. V.3 for the general construction). We can introduce the new manifold via a twisted product:
@@
The volume of the twisted manifold $\mathcal{M}^{3}_{g}$ scales as
$=$ $vol(\Sigma_{g}\times[0,1])=area(\Sigma_{g})=4\pi(g-1)=O(g),$ (89)
$I(\tau_{*}):=\{v\in H_{1}(\Sigma_{g};\mathbb{Z}_{2})\,|\,\tau_{*}(v)=v\}$ (90)
$\pi:\Sigma_{g}\to\Sigma_{g}/\langle\tau\rangle=:{{}_{g}S}$ (92)
[…]
In particular, the linear combination
$\tilde{\alpha}_{1}+\dots+\tilde{\alpha}_{N(\alpha)}=:v$ (96)
which implies
An easy application of Riemann-Hurwitz formula gives
$=$ $O(g/\log^{\frac{1}{2}}(g))=O(V/\log^{\frac{1}{2}}(V)),$ (100)
where the last equality is obtained via Eq. V.2.1.
One can prove the lower bound of the 1-systole $sys_{1}(\mathcal{M}^{3};\mathbb{Z}_{2})$$=$$\Omega(\log^{\frac{1}{2}}(V))$ by contradiction. We assume there exists a non-contractible loop $\gamma$ in $\mathcal{M}^{3}_{g}$ such that $length(\gamma)<O(\log^{\frac{1}{2}}(V))=O(\log^{\frac{1}{2}}(g))$. Consider lifting $\gamma$ in $\mathcal{M}^{3}_{g}$ to a curve $\tilde{\gamma}$ in the covering space $\Sigma_{g}\times\mathbb{R}$. The lifted curve $\tilde{\gamma}$ starts at point $(p,t)$ and ends at $(\tau^{m}p,t+m)$ for some $m\in\mathbb{Z}$. Since the bundle projection $\pi:\mathcal{M}^{3}_{g}\rightarrow[0,1]/\langle 0=1\rangle$ is length non-increasing, we have $|m|\leq length(\tilde{\gamma})=length(\gamma)<O(\log^{\frac{1}{2}}(g))$. Since we have $order(\tau)=\Omega(\log^{\frac{1}{2}}(g))$ from Eq. (85), we obtain $|m|<order(\tau)$, which suggests $p\neq\tau^{m}p$. Note that under the covering map $\sigma:\Sigma_{g}\rightarrow\Sigma_{g}/\langle\tau\rangle=:{{}_{g}S}$, both $p$ and $\tau^{m}p$ in $\Sigma_{g}$ are projected to the same point in ${}_{g}S$, therefore $p$ and $\tau^{m}p$ must differ by a non-trivial covering translation of the cover $\sigma$. However, any non-trivial covering translation must move each point of the total space $\Sigma_{g}$ by no less than twice the injectivity radius of the base ${}_{g}S$ since the geodisc balls of radius smaller than the injectivity radius $R(_{g}S)$ must be disjoint. According to Eq. (87), we have $R(_{g}S)\geq O(\log^{\frac{1}{2}}(g_{k}))$. Since the projection $\Sigma_{g}\times\mathbb{R}\rightarrow\Sigma_{g}$ is length nonincreasing, we have
[…]
So for some $t_{0}\in[0,1/2]$, we have $|z\cap\Sigma_{g}\times\{t_{0}\}|\leq 2Area(z)$. If $z\cap\Sigma_{g}\times\{t_{0}\}$ is non-trivial in $H_{1}(\Sigma_{g};\mathbb{Z}_{2})$ then from $sys_{1}(\Sigma_{g})\gtrsim log(g)$ we get that $Area(z)\gtrsim log(g)$. Otherwise, $z\cap\Sigma_{g}\times\{t_{0}\}$ bounds a 2-cycle $F\subset\Sigma_{g}\times\{t_{0}\}$. Using Buser’s inequality with $\lambda_{1}(\Sigma_{g})\gtrsim 1$, we can find a $F$ with $Area(F)\lesssim Area(z)$. But then the cycle $z+F-F$ can be lifted to $\Sigma_{g}\times\mathbb{R}$ from which we can see that $Area(z)+2Area(F)\geq Area(\Sigma_{g})\gtrsim g$. From the estimate on $Area(F)$ we get that $Area(z)\gtrsim g$. Thus, we see that $sys_{2}(\mathcal{M}_{g})\gtrsim\log(g)=\Omega(\log V)$. ∎
@@ V.2.2 Triple intersection and logical gates
[…]
@@
[…]
@@
There exists a family of fibre-bundle color code with encoding rate $O(1/\log^{\frac{1}{2}}(n))$ and distance lower bound $d$$=$$\Omega(\log^{\frac{1}{2}}(n))$, which supports fault-tolerant logical non-Clifford gates and parallelizable logical Clifford gates.
@@ V.3 Generic 3D hyperbolic codes via the mapping torus construction and constant-rate 3D hyperbolic codes
where for all $1\leq i\leq N\leq 2g$ the diagonal entries $A(i)$ are positive integers such that for all $i>1$, $A(i-1)$ divides $A(i)$. The diagonal matrix on the RHS is called the Smith normal form of $(\hat{f}-Id)$ and will be denoted as $SNF(\hat{f}-Id)$. The $SNF(\hat{f}-Id)$ does not depend on the pair $P,Q$ and is an invariant of the conjugacy class of $\hat{f}$ in $Sp(2g,\mathbb{Z})$. In particular,
@@
$1\displaystyle 1\leq Rank(H_{1}(\mathcal{M}(f);\mathbb{Z}))\leq 2g+$ 1 (121)
@@
$91.2439\displaystyle 23+16\sqrt{2}+4\sqrt{65+46\sqrt{2}}=$ 91.2439 (143)
@@
Figure 17: (a) Interaction graph for the logical CZ gates of the quasi-hyperbolic code. Colored edges with the same type (represented by both colors and line types) represent logical CZ gates that get simultaneously applied. Choosing certain vertices (represented by circles or ellipses) as logical qubits (L) and others as ancillae (A) set in the logical state |$\ket{0}$. (b) The effective interaction graph for the logical qubits (L). Note that each colored edge has a unique type, representing individually addressable and parallelizable logical CZ gate.
@@ VI.1 Generic scheme for parallelizable logical Clifford gates
Figure 19: (a) Interaction graph corresponding to logical CZ gates, where the edges represented by red dashed lines represent parallel logical CZ gates. One can selectively initialize ancillae in the second toric-code copy into $\overline{\ket{+}}$ state. (b) Circuits to apply parallel logical Hadamard via parallel logical CZ gates, logical-$X$ measurement, and conditional Z correction.
@@
The logical CZ gate structure for the quasi-hyperbolic code (as well as the other two codes up to a change of cycle labels) has been summarized in Eq. (69) and Eq. (70), and is also encoded in the interaction graph in Fig. 9(b). However, we see that there are two colored edges with the same type, meaning that a pair of logical CZ gates have to be applied simultanesouly and cannot be separated. In order to resolve this issue, we set a constant fraction of the logical qubits as ancilla qubits in the logical zero state |$\ket{\overline{0}}$ in order to let each logical CZ gate to be individually addressable.
As shown in Fig. 17(a), we choose a specific encoding scheme by encoding the logical information (labeled as $L$) into the logical qubit labeled as $(\Sigma_{g};2)$ (in the toric-code copy 2) while treating logical qubit labeled as $(\Sigma_{g};2)$ and $(\Sigma_{g};3)$ (in the toric-code copy 2 and 3) as ancilla (labeled as $A$). In the mean time, we encode logical information into logical qubit labeled as $(a^{(i)}\times c;1)$ and $(b^{(i)}\times c;1)$ in the toric-code copy 1, while setting the rest logical qubits in toric-code copy 2 and 3 as ancilla. Note that all the ancilla logical qubits ($A$) are fixed at the logical-0 state |$\ket{\overline{0}}$.
@@
Figure 24: Protocol to implement targeted logical CCZ gate between any selected triplet. (a) The steps of logical SWAP shuffle the selected triplet of logical qubits to the ancilla LDPC color code copy (equivalent to toric-code copy 4,5 and 6). (b) A logical CCZ gate is applied only to three selected logical qubits in toric-code copy 4, 5 and 6 since all the other logical qubits in the ancilla LDPC code copy are set in the |$\overline{\ket{0}}$ state.
@@
Since all the higher symmetry operator discussed in this paper are transversal logical gates, they them alone cannot be universal due to the Eastin-Knill theorem [74]. Therefore, we have to consider some other operations such as measurements. In particular, we now consider how to generate logical Hadamard gates with the logical CZ gates implemented by 1-form symmetries. As shown in Fig. 19(a), we can reinitialize a selected set of ancilla logical qubits in toric-code copy 2 in the logical state $\overline{\ket{+}}$ (+1 eigenstate of the logical-X operator). The reinitialization can be done by measuring logical-X membrane operators of these ancillas via a lattice surgery in the $O(1)$ neighborhood of a logical-X membrane [58]. After this re-initialization, a collection of colored edges (red-dashed lines) representing simultaneous logical CZ gates via transversal gate $\widetilde{\text{CZ}}^{(1,2)}_{\Sigma_{g}}$ applied on $\Sigma_{g}$ and between toric-code copy 1 and 2 are turned on. As shown in Fig. 19(b), one can apply these logical CZ gates in parallel and again measure the logical-X membranes corresponding of the corresponding logical qubits. After applying a conditional logical-Z correction via applying logical-Z strings, one effectively applies parallel logical Hadamard gates on all the selected logical qubits.
The parallelizability of such logical-$X$ measurements and hence the logical Hadamard gates are determined the packing properties of the thickened logical-$X$ membranes in the corresponding 3-manifolds, which will be studies in more details in future works. For the specific case of the quasi-hyperbolic code, the region of the lattice surgery are supported on a thickened torus $T^{2}\times[0,1]$. Its packing properties are determined by the packing of thickened loops $S^{1}\times[0,1]$ in each fibre, i.e., the hyperbolic surface $\Sigma_{g}$. Since the length of each loop is $O(\log(A))$ ($A$ being the area of the surface), we expect to pack at most $O(A/logA)=O(g/\log g)=O(k/\log(k))$ loops. Therefore, one obtains an upper bound of the parallelizability of the logical-$X$ meausrements and Hadamard gates as $O(k/\log(k))$. A lower bound will be obtained in future works by detailed study of the hyperbolic geometry.

### v3

@@ III.3 Logical CZ gates as 1-form symmetries
$U^{\prime}_{i,j}(\alpha_{2})=$ $\prod_{\beta^{1},\gamma^{1}}(-1)^{\int_{\mathcal{L}}(\hat{m}_{\beta}\beta^{1})\cup(\hat{l}_{\gamma}\gamma^{1})\cup\alpha^{1}}$
$=$ $\prod_{\beta^{1},\gamma^{1}}\left[(-1)^{\hat{m}_{\beta}\hat{l}_{\gamma}}\right]^{\int_{\mathcal{L}}\alpha^{1}\cup\beta^{1}\cup\gamma^{1}}$ (49)
$=$ $\prod_{\beta^{1},\gamma^{1}}\overline{\text{CZ}}[(\beta^{1};i),(\gamma^{1};j)]^{\int_{\mathcal{L}}{\alpha^{1}\cup\beta^{1}\cup\gamma^{1}}}$
$=$ $\prod_{\beta_{2},\gamma_{2}}\overline{\text{CZ}}[(\beta_{2};i),(\gamma_{2};j)]^{{|\alpha_{2}\cup\beta_{2}\cup\gamma_{2}|}},$ (50)
@@ IV Scaling of the code parameters for homological quantum codes defined on manifolds
In this section, we analyze the connection between the systolic geometry of a manifold and the code parameters of the corresponding homological qLDPC code. In particular, we show that the encoding rate and code distance corresponds to the Betti number and systoles of the underlying manifolds.
@@ IV.1 Review of hyperbolic geometry and codes
For DD==22, i.e., 2D arithemetic hyperbolic manifold $\mathcal{M}^{2}_{h}$, we can obtain the lower bound of the 1-systole
@@ IV.2 Betti number and encoding rate scaling in three and higher dimensions
where $\chi(\mathcal{M}^{2m})$ is the Euler characteristics and $Pf(\Omega)$ is the Pfaffian of the curvature form of the Levi-Civita connection. Since the Euler characteristics can be expressed as an alternating sum of the $i^{\text{th}}$ Betti number as $\chi(\mathcal{M}^{2m})=\sum_{i=0}^{2m}(-1)^{i}b_{i}$, it is in general difficult to obtain a simple scaling between the $i^{\text{th}}$ Betti number $b_{i}$ and the hyperbolic volume in dimension higher than two from this formula. The only exception is in 4D, where the 2nd Betti number is lower bounded by the volume in Ref.[58], leading to the corresponding (2,2)(2,2)-4D hyperbolic code 888(2,2) means both the logical-X and logical-Z operators are supported on 2-cycles. having constant encoding rate similar to the case of 2D hyperbolic code. Furthermore, in odd dimensions, the Chern-Gauss-Bonnet theorem simply does not apply.
@@
in a quite simple way from an information-theoretic perspective. For a given $D$-manifold, one can define D1D-1 copies of homological quantum codes (with $\lfloor D/2\rfloor$ distinct types due to Poincaré duality), where the physical qubits are assigned to the $i$-cells and the code spaces are associated with $\mathcal{H}_{(i)}=\mathbb{C}^{|H_{i}(\mathcal{M}^{D};\mathbb{Z}_{2})|}$ repsectively ($i=1,2,\cdots,D-1$). Consider the case that $\mathcal{M}^{D}$ only has one connected component, we have $b_{0}(\mathcal{M}^{D})=b_{D}(\mathcal{M}^{D})=1$. Therefore $\sum_{i=0}^{D}b_{i}(\mathcal{M}^{D})-2$ is just the total number of logical qubits in these D1D-1 copies of homological quantum codes. On the other hand, $vol(\mathcal{M}^{D})$ is just proportional to the total number of physical qubits. Therefore, we see that the asymptotic scaling of the Betti number has to be at most linear with the volume as suggested in Eq. (59), simply because one cannot encode more logical qubits than the number of physical qubits. Namely, a Betti number scaling like $O(vol(\mathcal{M}^{D})^{1+\epsilon})$ for any positive $\epsilon$ is impossible.
@@
Figure 8: (a) Illustration of the 3-torus $T^{3}$ as a product of a 2-torus $T^{2}$ and a circle $S^{1}$. (2) Three torus submanifold (2-cycles) $a\times c$, $b\times c$ and $a\times b$ have a triple intersection point at pp. (c) Illustration of the product manifold $\Sigma_{g}\times S^{1}$. (d) Submanifolds (2-cycles) $\Sigma_{g}$, $a^{(i)}\times c$ and $b^{(i)}\times c$ have a triple-intersection at point $p^{(i)}$. The total triple intersection number is $g$.
@@ V Code construction and logical gate structure for 3-manifolds
In this section, we investigate three classes of homological codes defined on 3-manifolds, which can all be viewed as a fibre-bundle construction or more specifically a mapping torus construction. We further investigate the triple intersection invariants and the corresponding logical CCZ and CZ gates in these codes.
In Sec. V.1, we first introduce the simplest construction: the quasi-hyperbolic code, which is formed by the product of a 2D hyperbolic code and a circle with logarithmic length. The code has almost-constant rate (up to logarithmic reduction) and logarithmic distance. This can be viewed as a fibre-bundle with a trivial (identity) twist. In Sec. V.2, we consider a fibre-bundle construction first proposed by Freedman, Meyer and Luo [59], where we take a twisted product of the 2D hyperbolic code and a circle with unit length, aiming to reduce the logarithmic overhead. We also study the triple intersection property in the presence of the twist. Nevertheless, we notice that the presence of the twist can reduce the homology which still leads to an almost-constant rate with a logarithmic reduction, while the distance also has a logarithmic lower bound. In Sec. V.3, we introduce a specific family of 3D hyperbolic codes called the Torelli mapping torus code, which are constructed from mapping tori (fibre bundle) of a pseudo-Anosov element in the Torelli subgroup. The pseudo-Anosov twist in this case does not lead to the reduction of homology and one hence maintain a constant rate, while the lower bound of the distance remains open.
@@ V.1.2 Triple intersection and logical gates
[…]
@@ V.2.1 Code construction and code parameters
We first review the construction of such 3-manifolds. We start with the choice of the fibre being a particular family of arithmetic hyperbolic 2-manifolds $\Sigma_{g}$ which are the quotient of the infinite hyperbolic plane $\mathbb{H}^{2}$ by a Fuchian group. We consider the co-compact torsion free Fuchscian group $\Gamma_{(-1,p)}$ derived from the type $\frac{-1,p}{\mathbb{Q}}$ quaternion algebra with the prime number p=3mod 4p=3\ $\text${mod}\ 4. One can explicitly write out the group $\Gamma_{(-1,p)}$ as
@@
    The map $\Sigma_{g}\rightarrow\Sigma_{g}/\langle\tau\rangle$ is a covering projection to the base surface $_{g}S=:\Sigma_{g}/\langle\tau\rangle$ (of genus <g<g), which has injectivity radius $R(_{g}S)\geq c_{3}(\log g)^{\frac{1}{2}}$.
where $c_{i}$ (ii=1,2,3) are positive constants independent of $g$. We now demonstrate the above properties in the following paragraphs.
@@
since both surfaces are obtained from compactifying $\mathbb{H}^{2}$ by quotienting out the corresponding isometry groups. Combined with the Gauss-Bonnet theorem Eq. (52), we know that the genus $g(N)\equiv genus(\mathbb{H}^{2}/\Gamma_{(-1,p)}(N))$ is proportional to the index $\Gamma_{(-1,p)}:\Gamma_{(-1,p)}(N)]$. Using Eq. (78), we obtain
@@
[…]
@@
In the above, we have demonstrated the properties of the hyperbolic surfaces $\Sigma_{g}$. We now construct a fibre bundle with $\Sigma_{g}$$\mathbb{H}^{2}/\Gamma_{(-1,p)}(N)$ as the fibre and a unit-length circle $S^{1}\equiv[0,1]$ as the basis, which is also called a mapping torus construction in this special case (see Sec. V.3 for the general construction). We can introduce the new manifold via a twisted product:
@@
See Appendix B.
@@ V.2.2 Triple intersection and logical gates
where the left hand side is the $\mathbb{Z}_{2}$ triple intersection number of the 2-cycles and $|\cdot\cap\cdot|$ on the right hand side is the $\mathbb{Z}_{2}$ intersection of two 1-cycles on $\Sigma_{g}$ [see Eq. (117) for precise definition].
@@
$\vline\,\cdot\cap\cdot|_{R}:I(\tau_{*})\times I(\tau_{*})$$\mathbb{Z}_{2}$$\vline\,\cdot\cap\cdot|_{\tau}:H_{1}(\Sigma_{g}/\tau;\mathbb{Z}_{2})\times H_{1}(\Sigma_{g}/\tau;\mathbb{Z}_{2})$$\mathbb{Z}_{2}$$\scriptstyle{Id}$ (95)
@@
[…]
@@ V.3 Generic 3D hyperbolic codes via the mapping torus construction and constant-rate 3D hyperbolic codes
where for all $1\leq i\leq N\leq 2g$ the diagonal entries $A(i)$ are positive integers such that for all $i>1$, A(i1)A(i-1) divides $A(i)$. The diagonal matrix on the RHS is called the Smith normal form of $(\hat{f}-Id)$ and will be denoted as $SNF(\hat{f}-Id)$. The $SNF(\hat{f}-Id)$ does not depend on the pair $P,Q$ and is an invariant of the conjugacy class of $\hat{f}$ in $Sp(2g,\mathbb{Z})$. In particular,
@@
$1\leq Rank(H_{1}(\mathcal{M}(f);\mathbb{Z}))\leq 2g+1$ (112)
@@
$23+16\sqrt{2}+4\sqrt{65+46\sqrt{2}}=91.2439$ (134)
@@
[…]
@@ VI.1 Generic scheme for parallelizable logical Clifford gates
[…]
@@
[…]
@@
[…]
@@
[…]
The parallelizability of such logical-$X$ measurements and hence the logical Hadamard gates are determined the packing properties of the thickened logical-$X$ membranes in the corresponding 3-manifolds, which will be studies in more details in future works. For the specific case of the quasi-hyperbolic code, the region of the lattice surgery are supported on a thickened torus $T^{2}\times[0,1]$. Its packing properties are determined by the packing of thickened loops $S^{1}\times[0,1]$ in each fibre, i.e., the hyperbolic surface $\Sigma_{g}$. Since the length of each loop is $O(\log(A))$ (AA being the area of the surface), we expect to pack at most $O(A/logA)=O(g/\log g)=O(k/\log(k))$ loops. Therefore, one obtains an upper bound of the parallelizability of the logical-$X$ meausrements and Hadamard gates as $O(k/\log(k))$. A lower bound will be obtained in future works by detailed study of the hyperbolic geometry.